OpenStudy (anonymous):

Help Please... You open a savings account and deposit $2,000. After 1 year of earning continuously compounded interest, your balance is $2,026.17. After 2 years, the balance is $2,052.68. Assuming you make no deposits or withdrawals, find the equation for the best-fitting exponential function to represent the balance of the account after x years. How much money will be in the account after 10 years?

OpenStudy (ankit042):

The problem starts by finding the interest rate 2000(1+r)=2026.17 r = 2026.17/2000-1 r = 13.09% Money after x years 2000(1+0.01309)^x

OpenStudy (anonymous):

OK

OpenStudy (ankit042):

Let me know if you need any explination

OpenStudy (anonymous):

One sec.

OpenStudy (anonymous):

So the answer would be 2277.66?

OpenStudy (ankit042):

Yeah something close to 2277